Skip to Search

 
#

Support Services Meeting Notes
December 3, 2007

Present: Bill Covey, Will Chaney, Kelley Cunningham, Brian Keith, Erich Kesse, Rob Roberts, Grace Strawn
Business Services
  1. Grace attended a meeting, given by the top people in Tigert, regarding the budget situation. It is looking dismal. Hopefully we will not have to give back money for next year's budget. UF is pushing for ways to save money. They are looking at positions. Apparently there are a lot of graphic designers at UF. They are considering consolidating graphic design services for campus as a whole. Raising tuition will continue. There is a cost reduction task force working on the issue. Priority is being given to increasing out-of-state grad student, but it is a bad situation the way things are currently structured. We have put forth nine proposals for next year's budget to the legislature. One proposals shows salary increases for faculty and staff. Exempt staff might be breaking the 40 hour work week. Outlook not so good.
  2. We've changed our reporting systems for OPS. It should be up and running this week.
Facilities
  1. Project in room 450 East is almost complete. There is one more section of shelving to move.
  2. IFAS material thing is hanging.
  3. Pete is back from Lutron training. It said it was good.
  4. Minor problems with locked doors in West over the weekend.
  5. Problem with elevator access to the 6th floor of West.
  6. Temperature problems in West.
  7. Security training for Facilities staff this Wednesday.
  8. Delay on card swipes - equipment problems and programming issues. Now the techs will be conducting the training, so it will be further delayed. Should be back on schedule next week. Hope to be off the fobs by the beginning of next year.
  9. Cleaning of the escalator - calls to the company have not been returned.
  10. Lighting in West - need to get the office switches working. The problem appears to be related to how they are grouped in the system.
Kelley
  1. Hikaru received both grants. One does not have to be set up in PeopleSoft, the other does.
  2. Chelsea Dinsmore applied for an ALA grant.
  3. John Ingram - cost share on the Baldwin grant - Grace is working on it.
Systems
  1. Backup Exec is up and running on the Exchange server. The continuous protection option will not work on a clustered Exchange server. We will be moving to campus AD soon enough.
  2. Public workstations to campus AD - we have tested adding some to AD. Xerox printing will work. In order to move all of the public machines, we will have to run a script utility. Amy is working on the script to move our domain to campus AD. We will let Public Services decide if they want to do the move all at once or group by group. Circa wants to test their new print solution on public machines on campus AD. We need to make sure that the public machines that switch over are not locked out from any database applications on our network that rely on our inherent sign-on. Will have to make staff aware that their staff log-on will no longer work on the public campus AD machines.
  3. 3480 tape conversion - all have been copied over to VTS. All that remains is to revisit a few problem tapes and verify the NL tapes that were converted.
  4. Blog/wiki server - operational and available for internal use. We are waiting for the Web 2.0 group to set policy. Marilyn asked up to set a wiki up right now, but we were unable as we want to limit it to Library staff. Need to set a policy. Tom Minton is working on a blog template. Blogs do not adhere to Web page policy.
DLC
  1. Laurie Taylor - RGP grant - CSE told her not to apply. Her research component is not strong enough. If this doesn't work out, there are two other places she can try to place the project.
  2. UFIR position posted.
  3. Brian's meeting with DLC went well.
  4. Effort Tracking forms - Erich wants to consult with Business Services on a few students.
  5. DLOC hired an incumbent that knows FIU both inside and out. FIU has resolved their internal problems.
  6. Mini-grant content is in - the student one is out. It may not be done until February.
  7. Bryant funds are sitting there - Erich may ask John Nemmers to get a temporary assignment to help administrate.
  8. DLC SAN - the server is set up and configured, but the backup is not yet configured. The procedures on portable disks are to be the same on the SAN. Will, Erich and Will to discuss this further after this meeting.
HR
  1. Deadline for Effort Tracking is tomorrow. We have received about 60% as of today. Will send out a deadline reminder.
  2. Despite the hiring freeze, we are recruiting for a number of positions - seven in all. 6.5 FTEs.
  3. Brian met with DLC.
  4. Faculty recruitment policy and procedures meeting with Denise Bennett scheduled this week. Need to update the policy.
  5. Received the T&P recommendations. Packets to go to DLC for scanning.
  6. Lane's time - termination not processed.
  7. Dec. 13 - Holiday Luncheon.
  8. Dec. 11 - Steve Fuquay's retirement.
  9. Revised org chart - need to follow-up with Trudi to figure out when we can get it up on the Web.
Directors
  1. Bill will be out of the office Wed - Fri afternoon.
  2. OPS hiring - need to make sure student time is entered before they leave for good. Need to delegate supervisors of OPS hires to make sure they get student time entered.
  3. Library Council was largely uneventful.

 

University of Florida Home Page